OpenStudy (anonymous):

Sam is observing the velocity of a car at different times. After two hours, the velocity of the car is 54 km/h. After four hours, the velocity of the car is 58 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used. (5 points) Part B: How can you graph the equation obtained in Part A for the first six hours?

OpenStudy (anonymous):

well i believe you have to find the avg acceleration first

ganeshie8 (ganeshie8):

say velocity = \(v\), and time = \(t\)

ganeshie8 (ganeshie8):

You are given : ` After two hours, the velocity of the car is 54 km/h. After four hours, the velocity of the car is 58 km/h.` that means (2, 54) and (4, 58) are two points on the required graph and u can use these to write the equation for velocity in slope-intercept form

ganeshie8 (ganeshie8):

knw how to write the equation of a line passing through two points ?

ganeshie8 (ganeshie8):

start by finding the slope

OpenStudy (anonymous):

(2,54) (4,58) 58-54=4/2= 2

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

Excellent ! so the required equation will be of form : \(v = 2t + b\) you need to find b by plugging in one of the given points

OpenStudy (anonymous):

Is it v=2t+50?

ganeshie8 (ganeshie8):

Perfect !

OpenStudy (anonymous):

Thanks how do I describe part B?

ganeshie8 (ganeshie8):

just find the `start point` and `end point` and connect them

OpenStudy (anonymous):

Would I plug six into the t?

ganeshie8 (ganeshie8):

Your equation for velocity from part A : \(v=2t+50\) when \(t=0,~~v = ?\) when \(t=6, ~~v=?\)

ganeshie8 (ganeshie8):

yes, plugging in t = 6 gives u the endpoint ^

OpenStudy (ageta):

When t=0 then v=50 and when t=6 then v=62